Twice a number equals 25 less than three times that same number. What is the number?
Answer by checkley77(12569)

2x=3x-25
2x-3x=-25
-x=-25
x=-25/-1
x=25 ans.
Proof:
2*25=3*25-254
50=75-25
50=50